What is molar conductivity and what is molar conductance ? What is the difference between both terms,?

Grade:12

1 Answers

Arun
6 years ago
The resistance of any conductor varies directly as its length (l) and inversely as its cross-sectional area (A), i.eMathematically  

R\ \alpha \ l

R\ \alpha \ \frac{1}{A}

\Rightarrow R = \rho \frac{l}{A}

Where ? is called the specific resistance.

l/A is known as cell constant.

If l = 1 cm and A = 1 cm2, then

R = ? 

The specific resistance is, thus defined as the resistance of one centimeter cube of a conductor.

The reciprocal of specific resistance is termed the specific conductance or it is the conductance of one centimeter cube of a conductor.

It is denoted by the symbo\kappa (kappa)l .

Thus, 

\kappa =\frac{1}{\rho }

Where\kappa =  kappa ?  the specific conductance             

Specific conductance is also called conductivity.

Further,

\rho =\frac{A\times R}{l}

\Rightarrow \frac{1}{\rho } =\frac{l}{A}\times \frac{1}{R}

\Rightarrow \kappa = \frac{l}{A}\times C

or  Specific conductance = Conductance × cell constant

In the case of electrolytic solutions, the specific conductance is defined as the conductance of a solution of definite dilution enclosed in a cell having two electrodes of unit area separated by one centimeter apart.

Unit of specific conductance: